# Defending Against Weight-Poisoning Backdoor Attacks for Parameter-Efficient Fine-Tuning

### v2

@@ 6.2 Results of Weight-Poisoning Backdoor Attack
We have also observed that P-tuning v2 exhibits lower ASR performance compared to P-tuning v1. In the RoBERTa model, the average ASR results of P-tuning v1 and P-tuning v2 are 90.43% vs. 58.83%. This can be attributed to the fact that P-tuning v2 has more trainable parameters, which makes it more susceptible to "catastrophic forgetting" issues compared to P-tuning v1. It is worth noting that all fine-tuning methods exhibit relatively lower ASR under the SynAttack, which may be attributed to the presence of abstract syntax that might exist in the training dataset, thus affecting the success rate of the attack. Nevertheless, the ASR of PEFT methods still surpasses that of full-tuning.
Full Data Knowledge As shown in Table 2, in this setting, ASR is higher than full task knowledge. For example, in the LLaMA model, the average ASR results of LoRA are 99.52% vs. 90.28%. Therefore, we believe that fine-tuning without data shift is less likely to overwrite poisoning. Similarly, the ASR of SynAttack is higher than full task knowledge. For experimental results pertaining to the CR and COLA datasets, please refer to Appendix B.
Hyperparameter Ablation Analysis Based on the analysis above, we found that the ASR degradation in PEFT is lower compared to the full-parameter fine-tuning method. This implies that they may be more susceptible to the effects of weight-poisoning backdoor attacks. Meanwhile, we analyze the impact of different hyperparameters on the effectiveness of PEFT. As depicted in Figs. 3(a), 3(b) and 3(c), the model exhibits a stable attack success rate as the virtual token and encoder hidden size increase. However, when faced with different learning rates, there are fluctuations in the standard deviation of the ASR. Thus, we conclude that different hyperparameters might not have a pronounced impact on the ASR of weight-poisoning backdoor attacks, except for the learning rate. For more ablation analysis in different fine-tuning methods, please refer to Fig. 4 in Appendix B.
@@
6.3 Results of Weight-Poisoning Attack Defense
We conducted a series of experiments to analyze and explain the effectiveness of our defense method under different settings. The baseline models include Back-translation (Back Tr.), ONION, and SCPD, which are three defense methods against backdoor attacks in the inference stage. Based on the results presented in Tables 1, 4, and 5 (Please see Appendix B), which are the full task knowledge setting, we can draw the following conclusions:

### v3

@@ 5.2 Results of Weight-Poisoning Backdoor Attack
We have also observed that P-tuning v2 exhibits lower ASR performance compared to P-tuning v1. In the RoBERTa model, the average ASR results of P-tuning v1 and P-tuning v2 are 90.43% vs. 58.83%. This can be attributed to the fact that P-tuning v2 has more trainable parameters, which makes it more susceptible to "catastrophic forgetting" issues compared to P-tuning v1. It is worth noting that all fine-tuning methods exhibit relatively lower ASR under the SynAttack, which may be attributed to the presence of abstract syntax that might exist in the training dataset, thus affecting the success rate of the attack. Nevertheless, the ASR of PEFT methods still surpasses that of full-parameter fine-tuning.
Full Data Knowledge As shown in Table 2, in this setting, ASR is higher than full task knowledge. For example, in the LLaMA model, the average ASR results of LoRA are 99.52% vs. 90.28%. Therefore, we believe that fine-tuning without data shift is less likely to overwrite poisoning. Similarly, the ASR of SynAttack is higher than full task knowledge. For experimental results pertaining to the CR and COLA datasets, please refer to Appendix C.
Hyperparameter Ablation Analysis Based on the analysis above, we found that the ASR degradation in PEFT is lower compared to the full-parameter fine-tuning method. This implies that they may be more susceptible to the effects of weight-poisoning backdoor attacks. Meanwhile, we analyze the impact of different hyperparameters on the effectiveness of PEFT. As depicted in Figs. 3(a), 3(b) and 3(c), the model exhibits a stable attack success rate as the virtual token and encoder hidden size increase. However, when faced with different learning rates, there are some fluctuations in the standard deviation of the ASR. Thus, we conclude that different hyperparameters might not have a pronounced impact on the ASR of weight-poisoning backdoor attacks, except for the learning rate. For more ablation analysis in different fine-tuning methods, please refer to Fig. 4 in Appendix C.
@@
5.3 Results of Weight-Poisoning Attack Defense
We conducted a series of experiments to analyze and explain the effectiveness of our defense method under different settings. The baseline models include Back-translation (Back Tr.), ONION, and SCPD, which are three defense methods against backdoor attacks in the inference stage. Based on the results presented in Tables 1, 4, and 5 (Please see Appendix C), which are the full task knowledge setting, we can draw the following conclusions: